# Diffusion Transformer Captures Spatial-Temporal Dependencies: A Theory for Gaussian Process Data

### v1

@@ B.3.2 Approximate Each Transformer Subnetworks
$\mathbf{s}_{i}^{(k+1)}$ $=\mathbf{s}_{i}^{(k)}-\eta_{t}\left[\sum_{j=1}^{N}(\alpha_{t}^{2}\bar{\bm{\Gamma}}_{ij}\bm{\Sigma})\mathbf{s}_{j}^{(k)}+\sigma^{2}_{t}\mathbf{s}_{i}^{(k)}-(\mathbf{x}_{i,t}-\alpha_{t}\bm{\mu}_{i})\right]$
$=\mathbf{s}_{i}^{(k)}-\underbrace{\eta_{t}\sum_{j=1}^{N}\alpha_{t}^{2}\bar{\bm{\Gamma}}_{ij}\bm{\Sigma}\mathbf{s}_{j}^{(k)}}_{(A)}+\underbrace{\left(-\eta_{t}\sigma_{t}^{2}\mathbf{s}_{i}^{(k)}+\eta_{t}(\mathbf{x}_{i,t}-\alpha_{t}\bm{\mu}_{i})\right)}_{(B)}.$ (13)
@@
$(A)=\eta_{t}\sum_{m=0}^{N-1}\sum_{j=1}^{N}\alpha^{2}\gamma_{m}\mathds{1}\{|i-j|=m\}\bm{\Sigma}\mathbf{s}_{j}.$
@@
$\qquad\qquad\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top}),f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\bigg{]}^{\top}.$
@@
$\qquad\qquad\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top}),f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\bigg{]}^{\top}.$
@@
$\qquad\qquad\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top}),f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\bigg{]}^{\top}.$
@@
$\mathbf{W}_{1}=\begin{bmatrix}-\mathbf{I}_{d}&\mathbf{0}_{d_{e}\times d_{e}}&\mathbf{0}_{d_{t}\times d_{t}}&\mathbf{0}_{d\times d}&\mathbf{0}_{d\times d}&\mathbf{I}_{d}&\mathbf{I}_{d}&\mathbf{I}_{d}&\mathbf{0}_{d\times(3d+1)}\\ \mathbf{I}_{d}&\mathbf{0}_{d_{e}\times d_{e}}&\mathbf{0}_{d_{t}\times d_{t}}&\mathbf{0}_{d\times d}&\mathbf{0}_{d\times d}&-\mathbf{I}_{d}&-\mathbf{I}_{d}&-\mathbf{I}_{d}&\mathbf{0}_{d\times(3d+1)}\\ &&&&\mathbf{I}_{d}&&&&\\ &&&&-\mathbf{I}_{d}&&&&\\ &&&&&\mathbf{I}_{d}&&&\\ &&&&&-\mathbf{I}_{d}&&&\\ &&&&&&\mathbf{I}_{d}&&\\ &&&&&&-\mathbf{I}_{d}&&\end{bmatrix}\in\mathbb{R}^{(8d)\times D},$
@@
$\begin{bmatrix}\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j})+f_{\rm mult}(\sigma^{2},\mathbf{s}_{i})-(\mathbf{x}_{i}-f_{\rm mult}(\alpha,\bm{\mu}_{i}))\\ -\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j})-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i})+(\mathbf{x}_{i}-f_{\rm mult}(\alpha,\bm{\mu}_{i}))\\ f_{\rm mult}(\alpha^{2},\mathbf{s}_{i})\\ -f_{\rm mult}(\alpha^{2},\mathbf{s}_{i})\\ f_{\rm mult}(\sigma^{2},\mathbf{s}_{i})\\ -f_{\rm mult}(\sigma^{2},\mathbf{s}_{i})\\ \sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j})\\ -\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j})\end{bmatrix}.$
@@
$=\Bigg{[}\mathbf{0}_{d}^{\top},\mathbf{0}_{d_{e}}^{\top},\mathbf{0}_{d_{t}}^{\top},-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top})-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i}^{\top})+(\mathbf{x}_{i}^{\top}-f_{\rm mult}(\alpha,\bm{\mu}_{i}^{\top})),$
$\qquad-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top}),-f_{\rm mult}(\alpha^{2},\mathbf{s}_{i}^{\top}),-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i}^{\top}),\mathbf{0}_{d}^{\top},0,\mathbf{0}_{3d}^{\top}\Bigg{]}^{\top}.$
@@
$\quad\mathbf{s}_{i}-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top})-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i}^{\top})+(\mathbf{x}_{i}^{\top}-f_{\rm mult}(\alpha,\bm{\mu}_{i}^{\top})),$
$\hskip 144.54pt\mathbf{0}_{d}^{\top},\mathbf{0}_{d}^{\top},\mathbf{0}_{d}^{\top},f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\Bigg{]}^{\top}.$
Finally, by implementing another multiplication module $f_{\rm mult}$ that scales the gradient by the learning rate $\eta_{t}$, we realize
$f_{\rm GD,2}$ $=f_{\rm mult}\circ{\mathcal{T}}\mathcal{B}_{2}\circ{\mathcal{T}}\mathcal{B}_{1}(\mathbf{y}_{i})=\Bigg{[}\mathbf{x}_{i}^{\top},\bm{e}_{i}^{\top},\bm{\phi}^{\top}(t),$
$\quad f_{\rm mult}\bigg{(}\eta_{t},\mathbf{s}_{i}-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top})-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i}^{\top})+(\mathbf{x}_{i}^{\top}-f_{\rm mult}(\alpha,\bm{\mu}_{i}^{\top}))\bigg{)},$
$\hskip 144.54pt\mathbf{0}_{d}^{\top},\mathbf{0}_{d}^{\top},\mathbf{0}_{d}^{\top},f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\Bigg{]}^{\top}.$
we have implemented one gradient descent iteration, where the output replaces $\mathbf{s}_{i}$ in the initial input vector. Consequently, we can repeat the multiplication module followed by the three transformer blocks to approximate the gradient descent for $K$ iterations as needed. Note that, we don’t need to calculate $f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})^{\top}$ after the first step of GD, which leads to the construction of $f_{\rm GD,2}$.
@@
$\widehat{\mathbf{s}}_{i}^{(k+1)}=\widehat{\mathbf{s}}_{i}^{(k)}-f_{\rm mult}\bigg{(}\eta_{t},\sum_{m=0}^{N-1}$ $\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\widehat{\mathbf{s}}_{j}^{(k)})$
$-f_{\rm mult}(\sigma^{2},\widehat{\mathbf{s}}_{i}^{(k)})+(\mathbf{x}_{i}^{\top}-f_{\rm mult}(\alpha,\bm{\mu}_{i}))\bigg{)}.$
@@
$\widehat{\mathbf{s}}_{i,\star}^{(k+1)}=\widehat{\mathbf{s}}_{i}^{(k)}-\eta_{t}\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot\alpha^{2}\widehat{\mathbf{s}}_{j}^{(k)}-\sigma^{2}\widehat{\mathbf{s}}_{i}^{(k)}+(\mathbf{x}_{i}^{\top}-\alpha\bm{\mu}_{i}).$
@@
We summarize our construction by characterizing the size of the architecture:
  • The input to the transformer is of dimension $D\times N$ with $D=9d+d_{e}+d_{t}+1$.
  • The number of transformer blocks is bounded by $L=2+L_{\rm mult}$. The number of attention heads in each block is bounded by $M=4J$. According to Corollary 3, the number of layers of the multiplication module $L_{\rm mult}=\mathcal{O}(\log{\left(\left\lVert\mathbf{x}\right\rVert_{\infty}\left\lVert\mathbf{s}\right\rVert_{\infty}Nd/\epsilon\right)})$.
  • By the construction above, the parameters in the transformer are bounded by $\max\{1,\frac{8}{\Delta}\eta\left\lVert\bm{\Sigma}\right\rVert_{\infty},r^{2}\}$. Here, $\left\lVert\bm{\Sigma}\right\rVert_{\infty}=\max_{ij}|\bm{\Sigma}_{ij}|$ denotes the largest magnitude of entries. Since there are at most $\mathcal{O}(d)$ nonzero weights in each weight matrix, we have the norm of parameter bounded by

### v2

@@ B.3.2 Approximate Each Transformer Subnetworks
$\mathbf{s}_{i}^{(k+1)}$ $=\mathbf{s}_{i}^{(k)}-\eta_{t}\left[\sum_{j=1}^{N}(\alpha_{t}^{2}\bar{\bm{\Gamma}}_{ij}\bm{\Sigma})\mathbf{s}_{j}^{(k)}+\sigma^{2}_{t}\mathbf{s}_{i}^{(k)}+(\mathbf{x}_{i,t}-\alpha_{t}\bm{\mu}_{i})\right]$
$=\mathbf{s}_{i}^{(k)}-\underbrace{\eta_{t}\sum_{j=1}^{N}\alpha_{t}^{2}\bar{\bm{\Gamma}}_{ij}\bm{\Sigma}\mathbf{s}_{j}^{(k)}}_{(A)}-\underbrace{\left(\eta_{t}\sigma_{t}^{2}\mathbf{s}_{i}^{(k)}+\eta_{t}(\mathbf{x}_{i,t}-\alpha_{t}\bm{\mu}_{i})\right)}_{(B)}.$ (13)
@@
$(A)=\eta\sum_{m=0}^{N-1}\sum_{j=1}^{N}\alpha^{2}\gamma_{m}\mathds{1}\{|i-j|=m\}\bm{\Sigma}\mathbf{s}_{j}.$
@@
$\qquad\qquad\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top}),f_{\rm mult}(\alpha,\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\bigg{]}^{\top}.$
@@
$\qquad\qquad\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top}),f_{\rm mult}(\alpha,\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\bigg{]}^{\top}.$
@@
$\qquad\qquad\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top}),f_{\rm mult}(\alpha,\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\bigg{]}^{\top}.$
@@
$\mathbf{W}_{1}=\begin{bmatrix}\mathbf{I}_{d}&\mathbf{0}_{d_{e}\times d_{e}}&\mathbf{0}_{d_{t}\times d_{t}}&\mathbf{I}_{d}&\mathbf{0}_{d\times d}&\mathbf{I}_{d}&\mathbf{I}_{d}&-\mathbf{I}_{d}&\mathbf{0}_{d\times(3d+1)}\\ -\mathbf{I}_{d}&\mathbf{0}_{d_{e}\times d_{e}}&\mathbf{0}_{d_{t}\times d_{t}}&-\mathbf{I}_{d}&\mathbf{0}_{d\times d}&-\mathbf{I}_{d}&-\mathbf{I}_{d}&\mathbf{I}_{d}&\mathbf{0}_{d\times(3d+1)}\\ &&&-\mathbf{I}_{d}&\mathbf{I}_{d}&&&&\\ &&&\mathbf{I}_{d}&-\mathbf{I}_{d}&&&&\\ &&&&&\mathbf{I}_{d}&&&\\ &&&&&-\mathbf{I}_{d}&&&\\ &&&&&&\mathbf{I}_{d}&&\\ &&&&&&-\mathbf{I}_{d}&&\end{bmatrix}\in\mathbb{R}^{(8d)\times D},$
@@
$\begin{bmatrix}\mathbf{s}_{i}+\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j})+f_{\rm mult}(\sigma^{2},\mathbf{s}_{i})+(\mathbf{x}_{i}-f_{\rm mult}(\alpha,\bm{\mu}_{i}))\\ -\mathbf{s}_{i}-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j})-f_{\rm mult}(\sigma^{2}_{t},\mathbf{s}_{i})-(\mathbf{x}_{i}-f_{\rm mult}(\alpha,\bm{\mu}_{i}))\\ -\mathbf{s}_{i}+f_{\rm mult}(\alpha^{2},\mathbf{s}_{i})\\ \mathbf{s}_{i}-f_{\rm mult}(\alpha^{2},\mathbf{s}_{i})\\ f_{\rm mult}(\sigma^{2},\mathbf{s}_{i})\\ -f_{\rm mult}(\sigma^{2},\mathbf{s}_{i})\\ \sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j})\\ -\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j})\end{bmatrix}.$
@@
$=\Bigg{[}\mathbf{0}_{d}^{\top},\mathbf{0}_{d_{e}}^{\top},\mathbf{0}_{d_{t}}^{\top},-\mathbf{s}_{i}^{\top}-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top})-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i}^{\top})-(\mathbf{x}_{i}^{\top}-f_{\rm mult}(\alpha,\bm{\mu}_{i}^{\top})),$
$\qquad\mathbf{s}_{i}^{\top}-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top}),-f_{\rm mult}(\alpha^{2},\mathbf{s}_{i}^{\top}),-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i}^{\top}),\mathbf{0}_{d}^{\top},0,\mathbf{0}_{3d}^{\top}\Bigg{]}^{\top}.$
@@
$\quad-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top})-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i}^{\top})-(\mathbf{x}_{i}^{\top}-f_{\rm mult}(\alpha,\bm{\mu}_{i}^{\top})),$
$\hskip 144.54pt\mathbf{s}_{i}^{\top},\mathbf{0}_{d}^{\top},\mathbf{0}_{d}^{\top},f_{\rm mult}(\alpha,\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\Bigg{]}^{\top}.$
Lastly, by implementing another multiplication module $f_{\rm mult}$ that scales the gradient by the learning rate $\eta_{t}$, we obtain
$f_{\rm mult}\circ{\mathcal{T}}\mathcal{B}_{2}\circ{\mathcal{T}}\mathcal{B}_{1}(\mathbf{y}_{i})=\Bigg{[}\mathbf{x}_{i}^{\top},\bm{e}_{i}^{\top},\bm{\phi}^{\top}(t),$
$\qquad f_{\rm mult}\bigg{(}\eta_{t},-\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\mathbf{s}_{j}^{\top})-f_{\rm mult}(\sigma^{2},\mathbf{s}_{i}^{\top})+(\mathbf{x}_{i}^{\top}-f_{\rm mult}(\alpha,\bm{\mu}_{i}^{\top}))\bigg{)},$
$\hskip 166.2212pt\mathbf{s}_{i}^{\top},\mathbf{0}_{d}^{\top},\mathbf{0}_{d}^{\top},f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})^{\top},1,\mathbf{0}_{3d}^{\top}\Bigg{]}^{\top}.$
It is straightforward to implement a feedforward layer to sum up $\mathbf{s}_{i}$ with the gradient increment in $f_{\rm mult}\circ{\mathcal{T}}\mathcal{B}_{2}\circ{\mathcal{T}}\mathcal{B}_{1}(\mathbf{y}_{i})$. Clearly, the feedforward layer can be realized by a transformer block with a trivial attention layer — similar to ${\mathcal{T}}\mathcal{B}_{2}$ with much simpler weight matrices. As a result, we have implemented one gradient descent iteration, where the output replaces $\mathbf{s}_{i}$ in the initial input vector. We denote $f_{\rm GD,2}$ as our network implementation of one GD iteration.
We can repeat the multiplication module followed by the two transformer blocks to approximate the gradient descent for $K$ iterations as needed. Note that, we don’t need to calculate $f_{\rm mult}(\alpha_{t},\bm{\mu}_{i})^{\top}$ after the first step of GD, which leads to the construction of $f_{\rm GD,2}$.
@@
$\widehat{\mathbf{s}}_{i}^{(k+1)}=\widehat{\mathbf{s}}_{i}^{(k)}+f_{\rm mult}\bigg{(}\eta_{t},-\sum_{m=0}^{N-1}$ $\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot f_{\rm mult}(\alpha^{2},\widehat{\mathbf{s}}_{j}^{(k)})$
$-f_{\rm mult}(\sigma^{2},\widehat{\mathbf{s}}_{i}^{(k)})-(\mathbf{x}_{i}^{\top}-f_{\rm mult}(\alpha,\bm{\mu}_{i}))\bigg{)}.$
@@
$\widehat{\mathbf{s}}_{i,\star}^{(k+1)}=\widehat{\mathbf{s}}_{i}^{(k)}-\eta_{t}\sum_{m=0}^{N-1}\sum_{j=1}^{N}\gamma_{m}\psi(\mathbf{e}_{i}^{\top}\mathbf{e}_{j})\bm{\Sigma}\cdot\alpha^{2}\widehat{\mathbf{s}}_{j}^{(k)}-\sigma^{2}\widehat{\mathbf{s}}_{i}^{(k)}-(\mathbf{x}_{i}^{\top}-\alpha\bm{\mu}_{i}).$
@@
Similar to the proof of Lemma 9, we summarize the network size for implementing one GD iteration (except the first iteration).
Table 2: Transformer size for approximating one GD iteration (except the first iteration)
Input dimension $D\times N$ with $D=9d+d_{e}+d_{t}+1$
# of blocks $L$ $2+L_{\rm mult}$ with $L_{\rm mult}=\mathcal{O}(\log{\left(\left\lVert\mathbf{x}\right\rVert_{\infty}\left\lVert\mathbf{s}\right\rVert_{\infty}Nd/\epsilon\right)})$
# of attention heads $M$ $4J$ with $J$ the covariance truncation length
Parameter bound $B$ $\mathcal{O}{\left(d(\left\lVert\bm{\Sigma}\right\rVert_{\infty}+r^{2}))+\left\lVert\mathbf{x}\right\rVert_{\infty}+\left\lVert\mathbf{s}\right\rVert_{\infty}\right)}$
To see the parameter bound $B$, we observe that the parameters in the constructed transformer are bounded by $\max\{1,\frac{8}{\Delta}\eta\left\lVert\bm{\Sigma}\right\rVert_{\infty},r^{2}\}$. Here, $\left\lVert\bm{\Sigma}\right\rVert_{\infty}=\max_{ij}|\bm{\Sigma}_{ij}|$ denotes the maximum magnitude of entries. Since there are at most $\mathcal{O}(d)$ nonzero weights in each weight matrix, we have the norm of the parameters bounded by